# Partial Label Learning with a Partner

### v1

@@ Abstract
In partial label learning (PLL), each instance is associated with a set of candidate labels among which only one is ground-truth. The majority of the existing works focuses on constructing robust classifiers to estimate the labeling confidence of candidate labels in order to identify the correct one. However, these methods usually struggle to rectify mislabeled samples. To help existing PLL methods identify and rectify mislabeled samples, in this paper, we introduce a novel partner classifier and propose a novel “mutual supervision” paradigm. Specifically, we instantiate the partner classifier predicated on the implicit fact that non-candidate labels of a sample should not be assigned to it, which is inherently accurate and has not been fully investigated in PLL. Furthermore, a novel collaborative term is formulated to link the base classifier and the partner one. During each stage of mutual supervision, both classifiers will blur each other’s predictions through a blurring mechanism to prevent overconfidence in a specific label. Extensive experiments demonstrate that the performance and disambiguation ability of several well-established stand-alone and deep-learning based PLL approaches can be significantly improved by coupling with this learning paradigm.
Introduction
As a representative framework of weakly supervised learning, partial label learning (PLL), where each instance is associated with a set of candidate labels among which only one is ground-truth, has garnered increasing attention in recent years (Cour, Sapp, and Taskar 2011; Han et al. 2018; Jin and Ghahramani 2002a; Papandreou et al. 2015; Ren et al. 2018; Zhou 2018; Zhu and Goldberg 2009; Chai, Tsang, and Chen 2020; Ren et al. 2018; Li, Guo, and Zhou 2021; Li and Liang 2019). PLL has also been applied to many real-world applications due to the growing demand for identifying the valid label from a set of candidate labels. For instance, in the automatic face naming task (Zeng et al. 2013; Guillaumin, Verbeek, and Schmid 2010), each face crapped from images or videos is associated with a list of names extracted from the corresponding title or caption (Gong, Yuan, and Bao 2022). Another example is the facial age estimation task: for each human face, the ages annotated by crowd-sourcing labelers are considered as candidate labels (Panis et al. 2016).
Formally speaking, suppose $\mathcal{X}=\mathbb{R}^{q}$ denotes the $q$-dimensional feature space and let $\mathcal{Y}=\{0,1\}^{l}$ be the label space with $l$ classes. Given a partial label data set $\mathcal{D}=\{\mathbf{x}_{i},S_{i}|1\leq i\leq n\}$ where $\mathbf{x}_{i}\in\mathcal{X}$, $S_{i}\subseteq\mathcal{Y}$ is the corresponding candidate label set and $n$ is the number of instances, the task of PLL is to induce a multi-class classifier $f:\mathcal{X}\rightarrow\mathcal{Y}$ based on $\mathcal{D}$. The most challenging part of PLL is that the ground-truth label $y_{i}$ of a training sample $\mathbf{x}_{i}$ conceals in its candidate labels, i.e, $y_{i}\in S_{i}$, which cannot be directly accessible during the training process.
@@ Introduction
To address this challenge, existing works focus on disambiguation, which involves differentiating the labeling confidences of each candidate label to identify the ground truth. This process typically relies on an alternative and iterative method for updating the classifier’s parameters. For instance, PL-AGGD (Wang, Zhang, and Li 2022) constructs a similarity graph to achieve disambiguation, and SURE (Feng and An 2019) aims to maximize an infinity norm to differentiate the ground-truth label. However, a significant yet rarely studied question arises in the context of such algorithms: can a classifier correct a false positive candidate label (i.e., invalid candidate label) with a large or upward-trending labeling confidence at a later stage? To explore this question, we conduct experiments on a real-world data set Lost (Cour et al. 2009) and record the labeling confidences of several candidate labels generated by PL-AGGD (Wang, Zhang, and Li 2022) in each iteration, which is shown in Fig. 1. Our findings reveal some intriguing phenomena:
@@
Figure 1: Two representative errors a typical PLL classifier, e.g., PL-AGGD (Wang, Zhang, and Li 2022) may make. $C_{GT}$ and $C_{GT}^{*}$ stand for the labeling confidence of the ground-truth label generated by PL-AGGD and PL-AGGD coupled with PLCP, and $C_{FP}$ and $C_{FP}^{*}$ stand for that of a false positive label predicted by PL-AGGD and PL-AGGD coupled with PLCP. (a). For a false positive candidate label with a large labeling confidence, although its confidence may decrease properly, it could still be larger than the ground-truth one’s. (b). The labeling confidence of a false positive candidate label keeps increasing and becomes the largest, which misleads the final prediction. When coupled with PLCP (vertical dashed line), the new labeling confidence of each candidate label generated by the partner classifier is adopted as the supervision to help PL-AGGD correct these errors, which results in a mutation in the figures.
@@
To address the aforementioned challenge, we propose a novel PLL mutual supervision framework called PLCP, referring to Partial Label Learning with a Classifier as Partner. Given a classifier provided by any existing PLL approach as the base classifier, as shown in Fig. 2, PLCP introduces an additional partner classifier to help the base classifier identify and rectify mislabeled samples. This partner classifier aims to provide more accurate and complementary information to the base classifier, thereby enabling better disambiguation and facilitating mutual supervision between the two classifiers. It is worth noting that the design of an effective partner classifier is crucial to the success of this framework, as the feedback from the partner classifier greatly influences the base classifier’s ability to identify and correct mislabeled samples. Since the information in non-candidate labels, which indicates that a set of labels DO NOT belong to a sample, is more precise than that in candidate labels and is often overlooked by the majority of existing works, the partner classifier can be designed to specify the labels that should not be assigned to a sample, thereby complementing the base classifier. Additionally, a novel collaborative term is also designed to link the base classifier and the partner classifier.
@@
In each stage of mutual supervision, the labeling confidence is first updated based on the base classifier’s modeling output. Subsequently, a blurring mechanism is employed to further process the labeling confidence to introduce uncertainty, which could potentially reduce the large confidence of some false positive candidate labels or increase the small confidence of the ground truth one. This updated labeling confidence then serves as the supervision information to interact with the partner classifier, whose final output will also be converted to supervise the base classifier. The predictions of the two classifiers, while distinct, are inextricably linked, enhancing the disambiguation ability of this paradigm in two opposing ways. With this mutual supervision paradigm, mislabeled instances have a higher likelihood of being corrected.
The main contributions of this paper can be summarized as follows:
  • We highlight two representative errors that a PLL classifier may make, which has not been previously investigated and offers a new insight into PLL.
  • We introduce a partner classifier based on non-candidate labels to better identify and correct mislabeled samples of a base classifier through a mutual supervision framework, which is applicable to all types of PLL approaches.
  • We propose a novel collaborative term in the partner classifier, which links the base classifier and itself. Additionally, a blurring mechanism is introduced to add uncertainty to the outputs, which effectively tackles the mentioned drawbacks.
  • We conduct experiments on several data sets to validate the effectiveness of this framework, and the results demonstrate that PLCP improves the disambiguation ability of the base classifier, leading to outstanding performance across all data sets.
@@
Related Work
Partial Label Learning
Partial label learning (PLL), also known as superset-label learning (Liu and Dietterich 2012, 2014) or ambiguous label learning (Hüllermeier and Beringer 2005; Zeng et al. 2013), is a representative weakly supervised learning framework which learns from inaccurate supervision information. In partial label learning, each instance is associated with a set of candidate labels with only one being ground-truth and others being false positive. As the ground-truth label of a sample conceals in the corresponding candidate label set, which can not be directly acquired during the training process, partial label learning task is a quite challenging problem.
To tackle the mentioned challenge, existing works mainly focus on disambiguation (Feng and An 2019; Nguyen and Caruana 2008; Zhang and Yu 2015; Wang, Zhang, and Li 2022; Fan et al. 2021; Xu, Lv, and Geng 2019; Zhang, Wu, and Bao 2022; Qian et al. 2023), which can be broadly divided into two categories: averaging-based approaches and identification-based approaches. For the averaging-based approaches (Hüllermeier and Beringer 2005; Cour, Sapp, and Taskar 2011; Zhang and Yu 2015), each candidate label of a training sample is treated equally as the ground-truth one and the final prediction is yielded by averaging the modeling outputs. For instance, PL-KNN (Hüllermeier and Beringer 2005) averages the candidate labels of neighboring samples to make the prediction. This kind of approach is intuitive, however, it can be easily influenced by false positive candidate labels which results in inferior performance. For identification-based approaches, (Feng and An 2018, 2019; Nguyen and Caruana 2008; Jin and Ghahramani 2002b; Yu and Zhang 2017), the ground-truth label is treated as a latent variable and can be identified through an iterative optimization procedure such as EM. Moreover, labeling confidence based strategy is proposed in many state-of-the-art identification based approaches for better disambiguation. (Zhang and Yu 2015) and (Wang, Zhang, and Li 2022) construct a similarity graph based on the feature space to generate labeling confidence of candidate labels.
Recently, deep-learning based models have been introduced to address PLL tasks (Lv et al. 2020; Xu et al. 2021; He et al. 2022; Wu, Wang, and Zhang 2022; Lyu, Wu, and Feng 2022; Xia et al. 2023). PICO (Wang et al. 2022) is a contrastive learning-based approach devised to tackle label ambiguity in partial label learning. This method seeks to discern the ground-truth label from the candidate set by utilizing contrastively learned embedding prototypes. Lv et al. proposed PRODEN in (Lv et al. 2020), a model where the simultaneous updating of the model and identification of true labels are seamlessly integrated. Furthermore, He et al. introduced a partial label learning method based on semantic label representations in (He et al. 2022). This method employs a novel weighted calibration rank loss to facilitate label disambiguation. By leveraging label confidence, the approach weights the similarity towards all candidate labels and subsequently yields a higher similarity of candidate labels in comparison to each non-candidate label.
However, as mentioned in Section 1, the above kinds of approach usually fail to identify and correct the mislabeled samples. to address this challenge, we propose a novel framework called PLCP in the next section.
Learning from Non-candidate Labels
The utilization of non-candidate labels (a.k.a. complementary labels) is not a novel concept. Earlier studies on complementary-label learning (CLL) (Ishida et al. 2017, 2019; Gao and Zhang 2021; Katsura and Uchida 2020) centered around classification tasks involving one single complementary label. In contrast to CLL, our approach involves the generation of complementary labels derived from non-candidate labels, serving the purpose of disambiguation in PLL.
The Proposed Approach
Denote $\mathbf{X}=[\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{n}]^{\mathsf{T}}\in\mathbb{R}^{n\times q}$ the sample matrix with $n$ instances, and $\mathbf{Y}=[\mathbf{y}_{1},\mathbf{y_{2}},...,\mathbf{y}_{n}]^{\mathsf{T}}\in\{0,1\}^{n\times l}$ the partial label matrix with $l$ labels, where $y_{ij}=1$ (resp. $y_{ij}=0$) if the $j$-th label of $\mathbf{x}_{i}$ resides in (resp. does not reside in) its candidate label set. Given the partial label data set $\mathcal{D}=\{\mathbf{x}_{i},S_{i}|1\leq i\leq n\}$, the task of PLL is to learn a multi-class classifier $f:\mathcal{X}\rightarrow\mathcal{Y}$ based on $\mathcal{D}$.
To address the challenge described in Section 1, a partner classifier $\mathcal{P}$ is designed to complement a base classifier $\mathcal{B}$ and also can be supervised by $\mathcal{B}$. $\mathcal{B}$ represents any existing PLL classifier with fine generality and flexibility. In each stage of mutual supervision, the base classifier updates the labeling confidence based on its modeling outputs, and then a blurring mechanism is applied to further process it. Subsequently the output of the labeling confidence is taken as the supervision information for the partner classifier. The learning pipeline of the partner classifier closely mirrors that of the base classifier. The following subsections will provide further details on this process.
Base Classifier
@@
Suppose $\mathbf{P}=[\mathbf{p}_{1},\mathbf{p}_{2},...,\mathbf{p}_{n}]^{\mathsf{T}}\in\mathbb{R}^{n\times l}$ is the labeling confidence matrix, where $\mathbf{p}_{i}\in\mathbb{R}^{l}$ represents the labeling confidence vector of $\mathbf{x}_{i}$, and $p_{ij}$ denotes the probability of the $j$-th label being the ground-truth label of $\mathbf{x}_{i}$. For the base classifier utilizing the labeling confidence strategy, $\mathbf{P}$ is initialized according to the base classifier. Otherwise, it is initialized as follows:
@@
The labeling confidence vector $\mathbf{p}_{i}$ typically satisfies the following constraints: $\sum_{j}p_{ij}=1$, $0\leq p_{ij}\leq y_{ij}$ (Wang, Zhang, and Li 2022; Feng and An 2018). The first constrains $\mathbf{p}_{i}$ to be normalized, and the second indicates that only the confidence of a candidate label has a chance to be positive. It should be noted that the ideal state of $\mathbf{p}_{i}$ is one-hot.
Once the base classifier has been trained, its modeling output will be generated. Denote the modeling output matrix $\mathbf{M}=[\mathbf{m}_{1},\mathbf{m}_{2},...,\mathbf{m}_{n}]^{\mathsf{T}}\in\mathbb{R}^{n\times l}$, where $\mathbf{m}_{ij}$ represents the probability of the $j$-th label being $\mathbf{x}_{i}$’s ground-truth label as predicted by the base classifier. It should be noted that $\mathbf{m}_{i}$ may either be a real-valued or one-hot vector depending on the specific base classifier. For instance, for SURE (Feng and An 2019) $\mathbf{m}_{i}$ is real-valued, while for PL-KNN (Hüllermeier and Beringer 2005) it is one-hot. Afterwards, $\mathbf{P}$ is updated to $\mathbf{P}_{1}$ through the following equation:
@@
Blurring Mechanism
In the next step, a blurring mechanism is designed to further process the labeling confidence matrix $\mathbf{P}_{1}$ to $\mathbf{Q}_{1}$ :
$\mathbf{Q}_{1}=\phi(e^{k}\mathbf{P}_{1})\odot\mathbf{Y},$ (3)
where $\phi(\cdot)$ is an element-wise operator, for a matrix $\mathbf{A}=[a_{ij}]_{n\times l}\in\mathbb{R}^{n\times l}$, $\phi(\mathbf{A})=[{\rm exp}(a_{ij})]_{n\times l}$. $k$ is a temperature parameter that controls the extent of blurring of labeling confidence. The labeling confidences are expected to be blurred at each stage of mutual supervision to prevent being overconfident in some false positive labels, hence we set $k<0$, which means two labeling confidences that differ significantly can also become close. $\odot$ represents the Hadamard product of two matrices. For matrices $\mathbf{A}$ and $\mathbf{B}$ with same size $n\times l$, $\mathbf{A}\odot\mathbf{B}=[a_{ij}b_{ij}]_{n\times l}$. The Hadamard product allows only the candidate label has a positive confidence. We then normalize each row of $\mathbf{Q}_{1}$ to satisfy the two constraints of labeling confidence, and output the result $\mathbf{O}_{1}\in\mathbb{R}^{n\times l}$.
Partner Classifier

### v3

# Appeal: Allow Mislabeled Samples the Chance to be Rectified in Partial Label Learning

@@ Abstract
In partial label learning (PLL), each instance is associated with a set of candidate labels among which only one is ground-truth. The majority of the existing works focuses on constructing robust classifiers to estimate the labeling confidence of candidate labels in order to identify the correct one. However, these methods usually struggle to identify and rectify mislabeled samples. To help these mislabeled samples “appeal” for themselves and help existing PLL methods identify and rectify mislabeled samples, in this paper, we propose the first appeal-based PLL framework. Specifically, we introduce a novel partner classifier and instantiate it predicated on the implicit fact that non-candidate labels of a sample should not be assigned to it, which is inherently accurate and has not been fully investigated in PLL. Furthermore, a novel collaborative term is formulated to link the base classifier and the partner one. During each stage of mutual supervision, both classifiers will blur each other’s predictions through a blurring mechanism to prevent overconfidence in a specific label. Extensive experiments demonstrate that the appeal and disambiguation ability of several well-established stand-alone and deep-learning based PLL approaches can be significantly improved by coupling with this learning paradigm.
Index Terms:
Partial label learning, appeal, disambiguation
1 Introduction
In 2002, Jin and Ghahramani [1] pioneered the concept of Partial Label Learning (PLL), a groundbreaking framework in which each instance is associated with multiple candidate labels, among which only one represents the actual ground-truth. Over the past two decades, the field of PLL has experienced remarkable growth [2, 3, 1, 4, 5, 6, 7, 8, 9, 10], driven by the increasing necessity to accurately discern the valid label from a set of potential candidates in diverse real-world applications. A notable example, as shown in Fig. 1, is the automatic face naming task [11, 12], where each facial image extracted from various media is linked to a list of names derived from corresponding titles or captions [13]. Another salient application is facial age estimation: for each human face, the ages annotated by crowd-sourcing labelers are considered as candidate labels [14].
Figure 1: Partial label learning in the automatic face naming task. The faces in this image can be automatically detected by a face detector, and then each face is assigned with the candidate name label set extracted from the script: Jennifer Aniston, Courteney Cox, Lisa Kudrow, Matthew Perry, David Schwimmer, and Matt LeBlanc, where only one is ground-truth.
@@ 1 Introduction
Evidently, the principal challenge in PLL lies in the inherent obscurity of the ground-truth label, which conceals within the candidate labels and remains inaccessible during the training phase. To surmount this obstacle, the most extensively researched and pivotal strategy in PLL-Disambiguation-emerges as critical. Existing methods predominantly achieve disambiguation by differentiating the labeling confidences of each candidate label to identify the ground truth. This process typically relies on an alternative and iterative optimization algorithm for updating the classifier’s parameters. For instance, LALO [15] implements constrained local consistency to differentiate the candidate labels, PL-AGGD [16] employs a similarity graph for effective disambiguation and PL-CL [17] adopts complementary information to help disambiguation. However, a significant yet rarely studied question comes to the front: can the instances that are incorrectly disambiguated, i.e., mislabeled, have the opportunity to be rectified? More specifically, can a classifier correct a false positive candidate label (i.e., invalid candidate label) with a large or upward-trending labeling confidence at a later stage? To explore this question, we conduct experiments on a real-world data set Lost [18] and record the labeling confidences of several candidate labels generated by PL-AGGD [16] in each iteration, which is shown in Fig. 2. Our findings reveal some intriguing phenomena:
@@
Figure 2: Two representative errors a typical PLL classifier, e.g., PL-AGGD [16] may make. $C_{GT}$ and $C_{GT}^{*}$ stand for the labeling confidence of the ground-truth label generated by PL-AGGD and PL-AGGD coupled with PLCP, and $C_{FP}$ and $C_{FP}^{*}$ stand for that of a false positive label predicted by PL-AGGD and PL-AGGD coupled with PLCP. (a). For a false positive candidate label with a large labeling confidence, although its confidence may decrease properly, it could still be larger than the ground-truth one’s. (b). The labeling confidence of a false positive candidate label keeps increasing and becomes the largest, which misleads the final prediction. When coupled with PLCP, the labeling confidence of each candidate label generated by the partner classifier is adopted as the supervision to help PL-AGGD correct these errors, which results in a mutation in the figures.
@@
We believe that these mislabeled samples should be afforded the opportunity to be corrected. In other words, these samples should have a chance to “appeal” on their behalf. In this case, the classifier’s disambiguation ability can also be further enhanced. Consequently, in this paper, we introduce another major strategy in PLL: Appeal, and propose the first appeal-based framework.
@@
Specifically, our framework PLCP, referring to Partial Label Learning with a Classifier as Partner, provides each instance with the chance to appeal. Given a classifier from any existed PLL approach as the base classifier, as shown in Fig. 3, PLCP integrates an additional partner classifier. This partner classifier assists the base classifier in identifying and rectifying mislabeled samples, offering more precise and complementary information to the base classifier, thereby enhancing disambiguation and fostering mutual supervision between the two classifiers. The design of the partner classifier is crucial in whether mislabeled samples could appeal for themselves, as the partner classifier’s feedback significantly influences the base classifier’s capability to recognize and amend mislabeled instances. Since the information in non-candidate labels, which indicates that a set of labels DO NOT belong to a sample, is typically more precise yet often overlooked by the majority of existing works, the partner classifier is devised as a complementary classifier. This classifier specifies the labels that should not be assigned to a sample, thereby complementing the base classifier. Additionally, a collaborative term is also designed to link the base classifier and the partner classifier.
@@
During mutual supervision, the labeling confidence is first updated based on the base classifier’s modeling output. A blurring mechanism is then applied to this updated labeling confidence, introducing uncertainty. This could potentially diminish the high confidence of certain false positive candidate labels or elevate the low confidence of the actual ground truth. This updated labeling confidence subsequently serves as the supervision information to interact with the partner classifier, whose final output, in turn, supervises the base classifier. The predictions of the two classifiers, while distinct, are inextricably linked, enhancing the disambiguation ability of this paradigm in two opposing ways. With this mutual supervision paradigm, the instances with disambiguation errors have a higher likelihood to appeal successfully.
The preliminary results of this work are presented in [19]. In this paper, we refine the method in [19] and propose a more robust version of PLCP. Besides, comprehensive theoretical analyses of PLCP are provided in this paper. The rest of the paper is organized as follows. We first introduce the appeal-based framework PLCP in Sec. 2, and then present the experimental results and ablation study in Sec. 3. Further analysis is provided in Sec. 4. We then review some related works in PLL and other related fields in Sec. 5. Finally, conclusion is given in Sec. 6
2 Appeal-based Framework: PLCP
Let $\mathcal{X}=\mathbb{R}^{q}$ denote the $q$-dimensional feature space and $\mathcal{Y}=\{0,1\}^{l}$ be the label space with $l$ classes. Given a partial label data set $\mathcal{D}=\{\mathbf{x}_{i},S_{i}|1\leq i\leq n\}$ where $\mathbf{x}_{i}\in\mathcal{X}$, $S_{i}\subseteq\mathcal{Y}$ is the corresponding candidate label set and $n$ is the number of instances. The task of PLL is to induce a multi-class classifier $f:\mathcal{X}\rightarrow\mathcal{Y}$ based on $\mathcal{D}$.
As shown in Fig. 3, a partner classifier $\mathcal{P}$ is designed to complement a base classifier $\mathcal{B}$ and also can be supervised by $\mathcal{B}$. $\mathcal{B}$ represents any existing PLL classifier with fine generality and flexibility. In each stage of mutual supervision, the base classifier updates the labeling confidence based on its modeling outputs, and then a blurring mechanism is applied to further process it. Subsequently the output of the labeling confidence is taken as the supervision information for the partner classifier. The learning pipeline of the partner classifier closely mirrors that of the base classifier. The following subsections will provide further details on this process.
@@ 2 Appeal-based Framework: PLCP
2.1 Base Classifier
@@
Denote $\mathbf{X}=[\mathbf{x}_{1},\mathbf{x}_{2},...,\mathbf{x}_{n}]^{\mathsf{T}}\in\mathbb{R}^{n\times q}$ the sample matrix with $n$ instances, and $\mathbf{Y}=[\mathbf{y}_{1},\mathbf{y_{2}},...,\mathbf{y}_{n}]^{\mathsf{T}}\in\{0,1\}^{n\times l}$ the partial label matrix, where $y_{ij}=1$ (resp. $y_{ij}=0$) if the $j$-th label of $\mathbf{x}_{i}$ resides in (resp. does not reside in) its candidate label set. Suppose $\mathbf{P}=[\mathbf{p}_{1},\mathbf{p}_{2},...,\mathbf{p}_{n}]^{\mathsf{T}}\in\mathbb{R}^{n\times l}$ is the labeling confidence matrix, where $\mathbf{p}_{i}\in\mathbb{R}^{l}$ represents the labeling confidence vector of $\mathbf{x}_{i}$, and $p_{ij}$ denotes the probability of the $j$-th label being the ground-truth label of $\mathbf{x}_{i}$. For the base classifier utilizing the labeling confidence strategy, $\mathbf{P}$ is initialized according to the base classifier. Generally, it is initialized as follows:
@@
The labeling confidence vector $\mathbf{p}_{i}$ typically satisfies the following constraints: $\sum_{j}p_{ij}=1$, $0\leq p_{ij}\leq y_{ij}$ [16, 15]. The first constrains $\mathbf{p}_{i}$ to be normalized, and the second indicates that only the confidence of a candidate label has a chance to be positive. It should be noted that the ideal state of $\mathbf{p}_{i}$ is one-hot.
Once the base classifier has been trained, its modeling output will be generated. Denote the modeling output matrix $\mathbf{M}=[\mathbf{m}_{1},\mathbf{m}_{2},...,\mathbf{m}_{n}]^{\mathsf{T}}\in\mathbb{R}^{n\times l}$, where $\mathbf{m}_{ij}$ represents the probability of the $j$-th label being $\mathbf{x}_{i}$’s ground-truth label as predicted by the base classifier. It should be noted that $\mathbf{m}_{i}$ may either be a real-valued or one-hot vector depending on the specific base classifier. For instance, for SURE [20] $\mathbf{m}_{i}$ is real-valued, while for PL-KNN [21] it is one-hot. Afterwards, $\mathbf{P}$ is updated to $\mathbf{P}_{1}$ through the following equation:
@@
2.2 Blurring Mechanism
In the next step, a blurring mechanism is designed to further process the labeling confidence matrix $\mathbf{P}_{1}$ to $\mathbf{O}_{1}\in\mathbb{R}^{n\times l}$ :
$\mathbf{Q}_{1}=$ $\mathcal{E}(e^{k}\mathbf{P}_{1})\odot\mathbf{Y},$ (3)
$[\mathbf{O}_{1}]_{i}$ $=\frac{[\mathbf{Q}_{1}]_{i}}{\|[\mathbf{Q}_{1}]_{i}\|},$
where $\mathcal{E}(\cdot)$ is an element-wise operator, for a matrix $\mathbf{A}=[a_{ij}]_{n\times l}\in\mathbb{R}^{n\times l}$, $\mathcal{E}(\mathbf{A})=[{\rm exp}(a_{ij})]_{n\times l}$. $k$ is a temperature parameter that controls the extent of blurring of labeling confidence. $\odot$ represents the Hadamard product of two matrices. For matrices $\mathbf{A}$ and $\mathbf{B}$ with same size $n\times l$, $\mathbf{A}\odot\mathbf{B}=[a_{ij}b_{ij}]_{n\times l}$. The Hadamard product allows only the candidate label has a positive confidence. We then normalize each row of $\mathbf{Q}_{1}$ and output the result $\mathbf{O}_{1}$.
The labeling confidences are expected to be blurred at each stage of mutual supervision to prevent being overconfident in some false positive labels, hence we set $k<0$ and then normalize the results, which means two labeling confidences that differ significantly can also become close. For more details on why the blurring mechanism can blur the outputs, please refer to Sec. 4.4.
2.3 Partner Classifier